OpenStudy (anonymous):

sove for b?? 3- (1/6)b=-3

OpenStudy (anonymous):

b=36

OpenStudy (anonymous):

\[3- (1/6)b=-3\] \[- (1/6)b=-6\] \[(1/6)b=6\] \[b=36\]